# COEVO: Co-Evolutionary Framework for Joint Functional Correctness and PPA Optimization in LLM-Based RTL Generation

### 16 Apr 2026 v1

@@ 1. Introduction
The growing complexity of modern integrated circuits has made manual Register-Transfer Level (RTL) design a costly bottleneck in hardware development (Liu et al., 2023a; Chang et al., 2023; Blocklove et al., 2023). Large language models (LLMs) (Yang et al., 2025b) offer a promising direction for automating RTL code generation from natural language specifications (Yang et al., 2025a; Xu et al., 2025), with standardized benchmarks such as VerilogEval (Liu et al., 2023b; Pinckney et al., 2025) and RTLLM (Liu et al., 2024c; Lu et al., 2024) accelerating progress. Extensive research has advanced functional correctness through domain-specific fine-tuning (Liu et al., 2024b; Pei et al., 2024; Zhao et al., 2025a; Cui et al., 2024), reinforcement learning with tool feedback (Teng et al., 2025; Wang et al., 2025a), and multi-agent frameworks (Liu et al., 2026; Zhao et al., 2025b; Ping et al., 2025; Abdollahi et al., 2026; Deng et al., 2026). However, functional correctness alone is insufficient for practical deployment, as LLM-generated RTL frequently exhibits suboptimal PPA compared to engineer-written implementations (Chen et al., 2025; Gubbi et al., 2025; Ping et al., 2026; Tasnia et al., 2025). Jointly optimizing functional correctness and PPA quality during spec-to-RTL generation has therefore become an important research direction.
@@ 2.1. LLM-Based RTL Code Generation
Automating RTL code generation from natural language specifications has attracted significant research attention. Domain-specific fine-tuning methods include RTLCoder (Liu et al., 2024b), BetterV (Pei et al., 2024), CodeV (Zhao et al., 2025a), OriGen (Cui et al., 2024), and ScaleRTL (Deng et al., 2025), while reinforcement learning methods such as VeriRL (Teng et al., 2025) and VeriReason (Wang et al., 2025a) incorporate compiler and simulator feedback into training. In parallel, agentic frameworks leverage general-purpose LLMs without modifying model parameters: MAGE (Zhao et al., 2025b) and VerilogCoder (Ho et al., 2025) decompose generation into specialized roles, VeriMoA (Ping et al., 2025) introduces quality-guided mixture-of-agents with multi-path generation, VeriSure (Liu et al., 2026) integrates formal verification with trace-driven temporal analysis, and ACE-RTL (Deng et al., 2026) combines an RTL-specialized LLM with a frontier reasoning model. These approaches have advanced functional correctness but do not systematically improve PPA quality.
@@ 5.1. Experimental Setup
Baselines. For functional correctness evaluation (RQ1), we compare against two categories of methods. Among agentic methods, we include I/O prompting, VeriOpt (Tasnia et al., 2025), VeriAgent (Wang et al., 2026), VerilogCoder (Ho et al., 2025), VeriMoA (Ping et al., 2025), REvolution (Min et al., 2026), and EvolVE (Hsin et al., 2026). Among training-based methods, we include SFT models (RTLCoder (Liu et al., 2024b), OriGen (Cui et al., 2024), HaVen (Yang et al., 2025c)) and SFT+RL models (VeriRL (Teng et al., 2025), QiMeng-SALV (Zhang et al., 2025), CodeV-R1 (Zhu et al., 2025)). For PPA comparison (RQ2), we compare against EvolVE, VeriAgent, and REvolution on RTLLM 2.0 using GPT-5.4-mini as the backbone.

### 17 Apr 2026 v2

@@ 1. Introduction
The growing complexity of modern integrated circuits has made manual Register-Transfer Level (RTL) design a costly bottleneck in hardware development (Liu et al., 2023a; Chang et al., 2023; Blocklove et al., 2023). Large language models (LLMs) (Yang et al., 2025b) offer a promising direction for automating RTL code generation from natural language specifications (Yang et al., 2025a; Xu et al., 2025), with standardized benchmarks such as VerilogEval (Liu et al., 2023b; Pinckney et al., 2025) and RTLLM (Liu et al., 2024c; Lu et al., 2024) accelerating progress. Extensive research has advanced functional correctness through domain-specific fine-tuning (Liu et al., 2024b; Pei et al., 2024; Zhao et al., 2025a; Cui et al., 2024), reinforcement learning with tool feedback (Teng et al., 2025; Wang et al., 2025a), and multi-agent frameworks (Ping et al., 2025b; Liu et al., 2026; Zhao et al., 2025b; Ping et al., 2025a; Abdollahi et al., 2026; Deng et al., 2026). However, functional correctness alone is insufficient for practical deployment, as LLM-generated RTL frequently exhibits suboptimal PPA compared to engineer-written implementations (Chen et al., 2025; Gubbi et al., 2025; Ping et al., 2026; Tasnia et al., 2025). Jointly optimizing functional correctness and PPA quality during spec-to-RTL generation has therefore become an important research direction.
@@ 2.1. LLM-Based RTL Code Generation
Automating RTL code generation from natural language specifications has attracted significant research attention. Domain-specific fine-tuning methods include RTLCoder (Liu et al., 2024b), BetterV (Pei et al., 2024), CodeV (Zhao et al., 2025a), OriGen (Cui et al., 2024), and ScaleRTL (Deng et al., 2025), while reinforcement learning methods such as VeriRL (Teng et al., 2025) and VeriReason (Wang et al., 2025a) incorporate compiler and simulator feedback into training. In parallel, agentic frameworks leverage general-purpose LLMs without modifying model parameters: MAGE (Zhao et al., 2025b) and VerilogCoder (Ho et al., 2025) decompose generation into specialized roles, VeriMoA (Ping et al., 2025a) introduces quality-guided mixture-of-agents with multi-path generation, VeriSure (Liu et al., 2026) integrates formal verification with trace-driven temporal analysis, and ACE-RTL (Deng et al., 2026) combines an RTL-specialized LLM with a frontier reasoning model. These approaches have advanced functional correctness but do not systematically improve PPA quality.
@@ 5.1. Experimental Setup
Baselines. For functional correctness evaluation (RQ1), we compare against two categories of methods. Among agentic methods, we include I/O prompting, VeriOpt (Tasnia et al., 2025), VeriAgent (Wang et al., 2026), VerilogCoder (Ho et al., 2025), VeriMoA (Ping et al., 2025a), REvolution (Min et al., 2026), and EvolVE (Hsin et al., 2026). Among training-based methods, we include SFT models (RTLCoder (Liu et al., 2024b), OriGen (Cui et al., 2024), HaVen (Yang et al., 2025c)) and SFT+RL models (VeriRL (Teng et al., 2025), QiMeng-SALV (Zhang et al., 2025), CodeV-R1 (Zhu et al., 2025)). For PPA comparison (RQ2), we compare against EvolVE, VeriAgent, and REvolution on RTLLM 2.0 using GPT-5.4-mini as the backbone.